# PrivaDE: Privacy-preserving Data Evaluation for Blockchain-based Data Marketplaces

### v1

@@ Abstract.
Evaluating the relevance of data is a critical task for model builders seeking to acquire datasets that enhance model performance. Ideally, such evaluation should allow the model builder to assess the utility of candidate data without exposing proprietary details of the model. At the same time, data providers must be assured that no information about their data—beyond the computed utility score—is disclosed to the model builder.
In this paper, we present PrivaDE, a cryptographic protocol for privacy-preserving utility scoring and selection of data for machine learning. While prior works (e.g., (Qian et al., 2024; Song et al., 2021)) have proposed data evaluation protocols, our approach advances the state of the art through a practical, blockchain-centric design. Leveraging the trustless nature of blockchains, PrivaDE enforces malicious-security guarantees and ensures strong privacy protection for both models and datasets. To achieve efficiency, we integrate several techniques—including model distillation, model splitting, and cut-and-choose zero-knowledge proofs—bringing the runtime to a practical level. Furthermore, we propose a unified utility scoring function that combines empirical loss, predictive entropy, and feature-space diversity, and that can be seamlessly integrated into active-learning workflows. Evaluation shows that PrivaDE performs data evaluation effectively, achieving online runtimes within 15 minutes even for models with millions of parameters.
@@
conference: ; ;
@@ 1. Introduction
As machine learning (ML) systems proliferate and generate substantial value, concerns have grown over large-scale, uncompensated data collection and the need to return value to original contributors. Collaborative learning frameworks (Sim et al., 2020), in which multiple parties pool data via a decentralized ledger, offer one remedy: trainers publish data requests, contributors opt in, and compensation for contributions is transparently recorded and fairly allocated. Despite the rise of blockchain-based data and ML marketplaces (e.g., (Rao, [n. d.])), rigorous mechanisms for data evaluation remain largely overlooked.
@@
The central challenge of data evaluation in collaborative learning is to select high-quality samples that maximally improve model performance while preserving the confidentiality of both data and model assets. Active learning (Dasgupta, 2011) addresses selection by querying the most informative examples—e.g., by uncertainty, expected loss, or diversity—thereby reducing annotation cost. Yet in a decentralized, privacy-sensitive marketplace, neither contributors nor model owners are willing to reveal raw features, labels, or model parameters prior to payment. It is therefore essential to design cryptographically secure protocols that compute data-utility scores under strong confidentiality guarantees and ensure correctness in a trustless environment.
@@
In this paper, we study how to design a secure protocol that enables a model owner (MO) and a data owner (DO) to jointly compute a data-utility score $\phi(M,D)$ for any model $M$ and dataset $D$, with practical applicability to blockchain-based data marketplaces. This score can subsequently serve as the basis for negotiation or automated payment for data acquisition in such marketplaces.
Our Contributions. To the best of our knowledge, prior work (e.g., (Song et al., 2021; Qian et al., 2024; Tian et al., 2022)) either operates in the semi-honest model or relies on a trusted third party (TTP) for most protocol steps. PrivaDE targets deployment in blockchain-based data marketplaces under realistic assumptions: we work in the malicious-adversary setting, where either the MO or the DO may deviate arbitrarily, and we require only minimal TTP involvement limited to setup and preprocessing. Both the model $M$ and the dataset $D$ are treated as private assets and are protected throughout the protocol.
PrivaDE considers the realistic setting where the DO holds a large dataset $D$, reflecting common scenarios in which a contributor possesses many samples. Our protocol enables the MO to obtain a dataset-level score $\phi(M,D)$ while minimizing passive information leakage by requiring only a representative subset for scoring. The scoring procedure is designed so that the subset-derived score is representative of the full dataset.
We introduce several design techniques (Section 4) that bring PrivaDE, under these strict assumptions, to practical runtimes. In particular, we combine model distillation, model splitting, and cut-and-choose ZKP to balance efficiency against potential privacy leakage.
@@
We further adapt PrivaDE to a blockchain-based data marketplace. Smart contracts deter misuse and enforce fairness in MPC by penalizing uncooperative parties via escrow slashing. The protocol can also be integrated with the data transaction to streamline the end-to-end marketplace workflow.
@@
Finally, we prove security in the simulation-based paradigm and demonstrate practicality through extensive experiments (Section 7) on real-world datasets and standard model architectures. Results show that PrivaDE can be flexibly tuned to achieve reasonable runtimes (as low as 45 seconds online for MNIST with LeNet), and that the scoring protocol is robust, effectively identifying high-performance datasets.
An overall design of PrivaDE can be seen in figure 1.
Figure 1. Overall Design of PrivaDE. Alice first prepares the model and Bob prepares the data, then they conduct secure inference together. The inference result is then used to decide a score on the dataset.
@@ 2.1. Machine learning
In machine learning, we compute a model $h$ from a training set $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $x_{i}$ has label $y_{i}$ (sometimes $z_{i}=(x_{i},y_{i})$). An algorithm $A$ searches a hypothesis space $\mathcal{H}$ for $h$ minimizing empirical loss $L_{D}(h)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(h(x_{i}),y_{i}\bigr)$, where $\ell$ quantifies the error between prediction $y^{\prime}=h(x)$ and true $y$. Common choices include squared loss $\ell(y^{\prime},y)=(y^{\prime}-y)^{2}$ for regression and zero-one loss $\ell(y^{\prime},y)=\mathbf{1}[y^{\prime}\neq y]$ for classification. For probabilistic outputs $\vec{y}^{\prime}$, cross-entropy $\ell_{\mathrm{CE}}(\vec{y}^{\prime},\vec{y})=-\sum_{i=1}^{d}\vec{y}^{i}\ln\vec{y}^{\prime i}$ is typical, where $\vec{y}$ is a one-hot vector. We denote models by $M$ and, when parameterized by $\theta$, by $M_{\theta}$.
@@ 2.2.1. Commitment schemes
We require two main security properties: hiding and binding. The hiding requires that no $\mathsf{PPT}$ adversary, upon receiving the output of $\mathtt{com}$, can figure out what message is committed in $\mathtt{com}$. Binding requires that $\mathtt{com}$ can be opened only to a specific message. We refer to App. B.1 for the formal definition of commitment.
@@ 2.2.2. Merkle Tree Commitments
The position-binding property (no two different $m$ open at the same $i$) holds under collision resistance of $H$. We refer to App. B.2 for the formal definition.
@@ 2.2.3. Zero-knowledge proofs
Informally, a NIZK prof is secure if it enjoys the properties of completeness, zero-knowledge and binding. Completeness simply requires that when the over and the verifier are honest the verifier always accepts the proof (i.e., it returns 1). Zero-knowledge requires that even when the verifier is malicious, the proof $\pi$ generated from a prover will leak no information about the witness $w$. Soundness requires that the verifier always rejects when receiving proofs for false statements. We refer to App. B.3 for the formal definition.
@@ 2.2.4. Multi-party computation
In short, a protocol is secure if a real-world adversary cannot gain an advantage over what they would learn in the ideal-world setting (that trivially captures the best possible security guarantee). When defining the security of a protocol $\Pi$, we will say that $\Pi$ securely realizes the function $f$, if $\Pi$ guarantees the above. In this paper we focus on the two-party setting, considering both security against semi-honest and malicious adversaries. In the former, the corrupted parties that attack the real-world protocol are assumed to follow the protocol description (but they can try to infer information from the honest parties’ inputs). In the malicious case, instead, parties can deviate arbitrarily from what the protocol prescribes. In App. B.4 we provide a formal definition of secure MPC and its variants.
@@ 2.3. Blockchain and Smart Contracts
A blockchain is a decentralized, append-only ledger that stores an ordered sequence of blocks; each block includes transactions, a timestamp, and the previous block’s hash, yielding immutability and tamper-evidence (Nakamoto, 2008; Bonneau et al., 2015). A smart contract is a program replicated and executed by blockchain nodes that enforces state transitions and agreements without intermediaries, enabling programmable, trustless transactions (Buterin, 2014; Wood, 2014).
@@
  • Enforce MPC fairness. Contracts escrow deposits and impose penalties for deviations, discouraging aborts and enforcing compliance during MPC.
@@
  • Support extensibility. The same infrastructure integrates with off-chain storage (e.g., IPFS) for availability and scalability.
@@ 3. Problem Description and Threat Model
We assume Bob contributes a larger dataset ($|D_{B}|>1$), reflecting practical scenarios such as a company representative contributing bulk records (e.g., sales data). To facilitate deployment on blockchain-based systems, PrivaDE is maliciously secure: it executes in the presence of a PPT adversary $\mathcal{A}$ who may arbitrarily deviate from the protocol, with the identity of the corrupted party fixed prior to execution. The protocol protects the honest party by detecting any deviation and safely aborting. PrivaDE requires no trusted party beyond the setup and preprocessing stages.
@@
4. Practical Optimizations for PrivaDE
In this section, we describe several design techniques that innovates PrivaDE and makes PrivaDE practical.
4.1. Model Distillation
Model distillation for modern neural networks, introduced by Hinton et al. (Hinton et al., 2015), is widely used to reduce model size and inference cost. The original (teacher) model guides training of a smaller (student) model, typically by matching the teacher’s logits.
Distillation is also common in secure inference (e.g., (Pereteanu et al., 2022)) to lower protocol runtime. Empirical studies (Romero et al., 2015; Cho and Hariharan, 2019) show that students with over $10\times$ fewer parameters can retain accuracy comparable to the teacher, depending on task and architecture.
In PrivaDE, we include an optional, local distillation step by the model owner prior to data evaluation. The distilled model is chosen to balance accuracy and size reduction. Beyond improving efficiency, distillation can also reduce leakage: due to limited capacity, a small student cannot faithfully encode all teacher information (Cho and Hariharan, 2019). We therefore deliberately select a student size that tolerates a minor accuracy loss in exchange for substantial runtime savings and reduced leakage. In Section 7, we show that a reduced model can still support effective data evaluation and produce results comparable to the original model.
4.2. Split Model
Most ML models, particularly neural networks, are compositions of layers: the output of one layer feeds the next until the final layer. This structure enables split learning (Vepakomma et al., 2018), which partitions a network at a cut layer. In a typical setup, the data owner trains the front segment and the model owner trains the back segment, coordinating forward and backward passes to reduce data exposure.
In secure inference, this idea yields split inference (e.g., (Pereteanu et al., 2022; Zhang et al., 2025a)): the front segment is executed within a secure protocol, while the remaining layers run in the clear. This reduces the cryptographic workload and thus latency, while limiting input leakage to the (carefully chosen) intermediate activations at the cut layer.
In PrivaDE, we also employ split-model inference, but propose a more sophisticated three-way split that factorizes $M$ as $C\circ B\circ A$ (Figure 2):
  1. (1)
    Block $A$. Contains the initial layers up to (and including) the first activation. We insert a secret mixer that permutes $A$’s outputs; this mixer can be an invertible linear layer or an invertible $1\times 1$ convolution that mixes channels (see (Kingma and Dhariwal, 2018)). Block $A$ is evaluated via a maliciously secure 2PC: Alice supplies weights, Bob supplies inputs. The output of $A$ is given only to Bob.
  2. (2)
    Block $B$. Comprises intermediate layers. The first layer of $B$ incorporates the inverse mixer to undo $A$’s permutation. The BBCC cut is chosen by the C2PI boundary-search procedure (Zhang et al., 2025a). Bob executes $B$ locally after receiving the corresponding weights from Alice.
  3. (3)
The BBCC split follows prior work aiming to curb input leakage while reducing runtime. We adopt the C2PI boundary search (Zhang et al., 2025a), which probes candidate cut layers by attempting input reconstruction from the cut-layer activations; for images, a split is accepted when the structural similarity index with the original input falls below $0.3$.
The AABB split further minimizes 2PC overhead by restricting cryptographic computation to the minimal front segment (up to the first activation), leaving the remainder in $B$. Because $B$’s weights are sent in the clear to Bob, $A$ prevents Bob from obtaining a self-sufficient front model that would yield reusable embeddings on arbitrary inputs.
Finally, we attach zero-knowledge proofs to the evaluations of $B$ and $C$ to enforce correctness and deter cheating. See Sec. 5 for protocol details.
4.2.1. Leakage of model $B$ to Bob:
Yokinski et al. (Yosinski et al., 2014) show that later layers tend to be more task-specific, whereas early layers capture generic features. Because neither the weights of $C$ nor the model outputs are revealed, Bob cannot reconstruct the full task model. However, $B\circ A$ may still yield useful embeddings and thus be valuable for fine-tuning on related domains; accordingly, we must bound the transferability of $B$.
To repurpose the model, Bob must:
  1. (1)
    Reconstruct $A$ to map raw inputs into $B$’s input space; and
  2. (2)
    Apply transfer learning to $B\circ A$ for other tasks.
For reconstructing $A$, Bob only observes his inputs and the corresponding outputs of $A$. We assume each contributor executes at most $k$ queries through $A$. A secret mixer (which can be different on different runs) applies a random invertible permutation to $A$’s outputs, which thwarts simple plug-and-play (attaching public pretrained fronts to $B$) and impedes collusion by multiple contributors pooling I/O pairs for extraction.
Empirical reports (Orekondy et al., 2019) suggest successful extraction often requires on the order of $10^{4}$ queries; thus, security improves when $k\ll 10^{4}$. To quantify robustness at budget $k$, we propose a reconstruction-under-budget experiment: randomly sample $\{x_{i}\}_{i=1}^{k}$ from $D_{A}$ and train $A^{\prime}$ (a re-initialized copy of $A$) by minimizing
with $\lambda>0$ weighting feature consistency through $B$. After training, fix $B$ and $C$, replace $A$ by $A^{\prime}$, and report top-1 accuracy on the standard test set. The primary metric is
A split is deemed robust if $\Delta\mathrm{Acc}$ is substantially negative at practical $k$, indicating that $A$ is not reconstructable and $B$ is not easily reusable under the stated constraints (cf. the privacy criterion in Zhang et al., 2025a).
We also assess the transferability of $B\circ A$ (hypothetically assuming $A$ is learned). Following You et al. (You et al., 2021), we compute LogME, the log marginal evidence that a Bayesian linear regressor explains labels from features. Concretely, we pass labeled images through $B(A(x))$ to obtain features and take the resulting LogME value as a transferability score; lower scores indicate poorer linear transfer and thus reduced leakage risk.
Alice decides by herself if the leakage is acceptable based on $\Delta\mathrm{Acc}$ and the LogME score by comparing it to baselines of other models in the marketplace.
4.2.2. Leakage of activations $B(A(x))$ to Alice:
We quantify input leakage by reconstruction quality, training an inverse model that maps $B(A(x))$ back to $x$. For images, we adopt DINA (Zhang et al., 2025a) and report the average structural similarity index (SSIM) over a test set; we accept the split when the mean SSIM <0.3<0.3, indicating low-fidelity reconstructions.
4.3. Representative Set Selection
To mitigate this, our protocol requires Bob to extract a much smaller representative subset $D_{R}\subset D_{B}$, and then jointly and securely evaluate Alice’s model only on $D_{R}$. To ensure correctness even if Bob is malicious, we enforce honest sampling via a cut-and-choose subroutine (Sec. 4.3.2). Similar representative-sampling strategies (without security) are standard in batch active learning (Sener and Savarese, 2018) and related ML tasks (Mirzasoleiman et al., 2013).
We formalize representativeness as follows:
Definition 4.1 (Representativeness of $D_{R}$ to $D_{B}$).
Let $D_{R}\subset D_{B}$, and fix a distance threshold $d>0$ and outlier tolerance $\delta\in[0,1]$. Then $D_{R}$ is $(d,\delta)$-representative if $\Pr_{x\sim D_{B}}\!\bigl[\exists\,x^{\prime}\in D_{R}:\|x-x^{\prime}\|<d\bigr]\geq 1-\delta$, where the probability is with respect to the empirical distribution over $D_{B}$.
We write $\mathcal{R}(D_{B})$ for any algorithm that, given $|D_{B}|=n>k$, returns $D_{R}\subset D_{B}$ of size $k$. Common methods include $k$-center clustering and submodular maximization (Nemhauser et al., 1978; Mirzasoleiman et al., 2013). The parameters $(k,d,\delta)$ are chosen jointly by Alice and Bob per application.
4.3.1. Dimension Reduction.
The selection step involves pairwise distances as in Def. 4.1, which becomes expensive in $\mathbb{R}^{d}$ as $d$ grows. We therefore apply random linear projections (Bingham and Mannila, 2001):
$x^{\prime}=R\,x,\qquad R\in\mathbb{R}^{m\times d},\;m\ll d,$
which preserve distances up to $\varepsilon$-distortion by the Johnson–Lindenstrauss lemma (Johnson and Lindenstrauss, 1984). In a secure setting, Alice and Bob jointly sample $R$ via $md$ secure coin flips (Blum, 1981), deriving a shared pseudorandom seed. Bob projects his data locally and runs the representative-selection protocol on $\{x^{\prime}\}$, reducing computation and communication while retaining scoring fidelity.
4.3.2. Challenge Protocol for Representative Set Selection
A naive way to securely implement the representative-set algorithm $\mathcal{R}$ is via a maliciously secure two-party computation (2PC). However, because state-of-the-art clustering often relies on costly Euclidean distances, this approach incurs substantial overhead.
As an alternative, we introduce the challenge protocol $\mathsf{CP}$. Bob commits each point $x_{i}\in D_{B}^{x}$ ($D_{B}^{x}:=\{x\mid(x,y)\in D_{B}\}$) via commitments $\mathtt{com}_{1},\dots,\mathtt{com}_{n}$, and publishes the representative set $D_{R}$ by revealing its indices $I_{R}:=\{i|x_{i}\in D_{R}^{x}\}$.
Alice selects a random index $i\in[n]$ and challenges Bob to prove, in zero knowledge, that the committed point $x_{i}$ is well represented by $D_{R}$. Bob recovers $x_{i}$, computes $d_{i}^{\prime}=\min_{x^{r}_{j}\in D_{R}^{x}}\|x^{r}_{j}-x_{i}\|$, and:
  • otherwise replies $\mathsf{fail}_{i}$.
By repeating this challenge, Alice gains confidence that $D_{R}$ is representative of $D_{B}$. The full protocol appears in Fig. 3, and its security is established in the next section.
If Alice requires $D_{R}$ to be $(d,\delta)$-representative of $D_{B}$, then running $|I|=\Theta(\ln n/\delta)$ challenges suffices to ensure, with probability at least $1-n^{-c}$, that at most a $\delta$ fraction of points are outliers after $c\ln n/\delta$ successful proofs. In the full PrivaDE protocol, $d$ and $\delta$ are pre-agreed, and $I\subset[n]$ denotes the indices Alice selects for $\mathsf{CP}$.
4.4. Audit-Style Cut-and-Choose for ML Inference
While significant research has focused on zero-knowledge proofs (ZKPs) for verifying ML inference (e.g., (Huang et al., 2022; Kang et al., 2022; Liu et al., 2021)), constructing end-to-end proofs remains memory- and computation-intensive.
Instead of requiring a full proof, we propose a lightweight protocol that detects incorrect batched inference with tunable detection probability (e.g., 80–95%) by sampling and proving only a small, random subset of local layer transitions. The protocol $\mathsf{CnCZK}$ serves as a drop-in replacement for a NIZKPoK of the following language $L_{\textsf{ML-Inf}}$. We consider two scenarios: (1) the model weights and outputs are hidden as part of the witness and the data point is public (blue); (2) the data point is hidden as part of the witness and the model weights and outputs are public (red). For each scenario, read the black text together with the text in the corresponding colour.
Notation.
The protocol $\mathsf{CnCZK}$ proceeds as follows.
4.4.1. Cheating-detection probability
Theorem 4.2.
The proof appears in Appendix A.1. As an illustration, with $N=100$, $L=10$, $\rho=0.10$, $m=25$, and $s=6$, the detection probability exceeds $80\%$. This setting audits only $ms/(NL)=150/1000=15\%$ of all layer evaluations, yielding about an $85\%$ reduction relative to checking every layer of every point.

### v2

@@ Abstract.
Evaluating the usefulness of data before purchase is essential when obtaining data for high-quality machine learning models, yet both model builders and data providers are often unwilling to reveal their proprietary assets.
We present PrivaDE, a privacy-preserving protocol that allows a model owner and a data owner to jointly compute a utility score for a candidate dataset without fully exposing model parameters, raw features, or labels. PrivaDE provides strong security against malicious behavior and can be integrated into blockchain-based marketplaces, where smart contracts enforce fair execution and payment. To make the protocol practical, we propose optimizations to enable efficient secure model inference, and a model-agnostic scoring method that uses only a small, representative subset of the data while still reflecting its impact on downstream training. Evaluation shows that PrivaDE performs data evaluation effectively, achieving online runtimes within 15 minutes even for models with millions of parameters.
@@
conference: ACM ASIA Conference on Computer and Communications Security; June 01–06, 2026; Bangalore, India
@@ 1. Introduction
As machine learning (ML) systems proliferate and generate substantial value, concerns have grown over large-scale, uncompensated data collection and the need to return value to original contributors. Collaborative learning frameworks (collabML), in which multiple parties pool data via a decentralized ledger, offer one remedy: trainers publish data requests, contributors opt in, and contributions are compensated in a transparent and fair manner. Despite the rise of blockchain-based data and ML marketplaces (e.g., (rao_bittensor_whitepaper)), rigorous mechanisms for data evaluation remain largely overlooked.
@@
The central challenge of data evaluation in collaborative learning is how to select high-quality samples that most improve model performance while preserving the confidentiality of both data and model assets. The problem is magnified in decentralized, privacy-sensitive marketplaces, where neither contributors nor model owners are willing to reveal raw features, labels, or model parameters prior to payment. It is therefore essential to design cryptographically secure protocols that compute data-utility scores under strong confidentiality guarantees while ensuring correctness in a trustless environment.
@@
In this paper, we design a secure protocol that enables a model owner (MO) and a data owner (DO) to jointly compute a data-utility score $\phi(M,D)$ for a model $M$ and dataset $D$, with practical applicability to blockchain-based data marketplaces. This score can subsequently serve as the basis for negotiation or automated payment for data acquisition in such marketplaces.
Our Contributions. We introduce PrivaDE, a two-party, maliciously secure protocol for dataset scoring. PrivaDE has two main components: (i) an improved and more efficient secure model inference protocol that obtains the model’s predictions on the dataset, and (ii) a data scoring protocol that takes the data points and model predictions as input and outputs a score $\phi$ representing the usefulness of the data.
Secure model inference under malicious security, e.g., with SPDZ (mpspdz) or MASCOT (mascot), is known to incur very high overhead. We introduce three optimization techniques (Section 5) that make the secure inference component of PrivaDE practical under a strict malicious-adversary assumption. In particular, we combine model distillation (to shrink the model), model splitting (to reduce the portion of computation run in secure environments), and a cut-and-choose ZKP audit (to reduce verification cost) to carefully balance efficiency and security.
We further design a novel, model-agnostic scoring function (Alg. 1) for the data scoring component, which evaluates data points and yields a score that correlates with downstream training performance. The scoring function builds on active learning criteria, namely diversity (how spread out the points are) and uncertainty (how uncertain the model’s predictions are), which have been empirically shown to be effective heuristics for data quality. Our scoring function is easy to compute under maliciously secure protocols and minimizes passive information leakage by operating on a representative subset of the dataset. The scoring procedure is constructed so that the subset-derived score accurately reflects the full dataset.
@@
We provide a brief proof-of-concept integration of PrivaDE into a blockchain-based data marketplace. Smart contracts deter misuse and enforce fairness in MPC by penalizing uncooperative parties via escrow slashing, and the protocol can be tightly coupled with the data transaction to streamline the end-to-end marketplace workflow.
@@
Finally, we prove security in the simulation-based paradigm and demonstrate practicality through extensive experiments (Section 7) on real-world datasets and standard model architectures. Our results show that PrivaDE can be flexibly tuned to achieve reasonable runtimes (as low as 45 seconds of online time for MNIST with LeNet), and that the scoring protocol is robust, effectively identifying high-value datasets.
To our knowledge, this is the first work to propose a secure data evaluation scheme in the malicious setting together with a model-agnostic, crypto-friendly scoring function. Prior work (e.g., (songDE; qianDE; privatedv)) either operates in the semi-honest model or relies on a trusted third party (TTP) for most protocol steps. An overview of PrivaDE is shown in Fig. 1.
Figure 1. Overall design of PrivaDE. The core protocol consists of two parts: a secure inference of the data points and a scoring function that takes the inference results and labeled data points as input. Several optimization components, detailed in Section 5, are applied before secure inference to reduce the overhead of the first component.
Organization.
Section 2 reviews background on machine learning, cryptography, and blockchain. Section 3 formalizes the problem and threat model. Section 4 presents the full protocol and security proofs. Section 5 explains optimization techniques applied that reduce the overhead of secure inference. Section 6 describes a proof-of-concept integration of PrivaDE into a blockchain-based marketplace, Section 7 evaluates the practicality and robustness of our construction, and Section 8 discusses related work.
@@ 2.1. Machine learning
In supervised machine learning, we compute a model $h$ from a training set $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $x_{i}$ has label $y_{i}$ (sometimes $z_{i}=(x_{i},y_{i})$). An algorithm $A$ searches a hypothesis space $\mathcal{H}$ for $h$ minimizing empirical loss $L_{D}(h)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(h(x_{i}),y_{i}\bigr)$, where $\ell$ quantifies the error between prediction $y^{\prime}=h(x)$ and true $y$. Common choices include squared loss $\ell(y^{\prime},y)=(y^{\prime}-y)^{2}$ for regression and zero-one loss $\ell(y^{\prime},y)=\mathbf{1}[y^{\prime}\neq y]$ for classification. For probabilistic outputs $\vec{y}^{\prime}$, cross-entropy $\ell_{\mathrm{CE}}(\vec{y}^{\prime},\vec{y})=-\sum_{i=1}^{d}\vec{y}^{i}\ln\vec{y}^{\prime i}$ is typical, where $\vec{y}$ is a one-hot vector. We denote models by $M$ and, when parameterized by $\theta$, by $M_{\theta}$.
@@ 2.2.1. Commitment schemes
We require two main security properties: hiding and binding. Hiding ensures no $\mathsf{PPT}$ adversary can determine the committed message from $\mathtt{com}$. Binding ensures $\mathtt{com}$ can be opened to only one specific message. See App. C.1 for formal definitions.
@@ 2.2.2. Merkle Tree Commitments
The position-binding property (no two different $m$ open at the same $i$) holds under collision resistance of $H$. We refer to App. C.2 for the formal definition.
@@ 2.2.3. Zero-knowledge proofs
Informally, a NIZK proof is secure if it satisfies completeness, zero-knowledge, and soundness. Completeness requires honest verifiers always accept honest proofs. Zero-knowledge ensures proofs leak no information about the witness $w$, even to malicious verifiers. Soundness requires verifiers reject proofs of false statements. We refer to App. C.3 for the formal definition.
@@ 2.2.4. Multi-party computation
In short, a protocol is secure if a real-world adversary cannot gain an advantage over what they would learn in the ideal-world setting (that trivially captures the best possible security guarantee). When defining the security of a protocol $\Pi$, we will say that $\Pi$ securely realizes the function $f$, if $\Pi$ guarantees the above. In this paper we focus on the two-party setting, considering security against malicious adversaries. Parties can deviate arbitrarily from what the protocol prescribes. In App. C.4 we provide a formal definition of secure MPC and its variants.
@@ 2.3. Blockchain and Smart Contracts
A blockchain is a decentralized, append-only ledger storing an ordered sequence of blocks; each block includes transactions, a timestamp, and the previous block’s hash, yielding immutability (nakamoto2008bitcoin; bonneau2015sok). A smart contract is a program replicated and executed by blockchain nodes, enforcing state transitions without intermediaries (buterin2014next; wood2014ethereum).
@@
  • Enforce MPC fairness. Contracts escrow deposits and impose penalties for deviations, discouraging aborts and enforcing compliance during MPC (fairmpc).
@@ 3. Problem Description and Threat Model
We assume Bob contributes a dataset with multiple points ($|D_{B}|>1$), reflecting practical scenarios such as a company representative contributing bulk records (e.g., sales data). To facilitate deployment on blockchain-based systems, PrivaDE is maliciously secure: it executes in the presence of a PPT adversary $\mathcal{A}$ who may arbitrarily deviate from the protocol, with the identity of the corrupted party fixed prior to execution. The protocol protects the honest party by detecting any deviation and safely aborting. PrivaDE requires no trusted party beyond the setup and preprocessing stages.